Dear Richard dear all,

please find below a short summary of the activities of today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni, M. Xompero
ADS: P. Lazzarini
LBTO: G. Brusa

Total # of person: 9

W unit

1) Functionality test of the unit was repeated successfully.
2) temperature control of the CCD39 electronic inside the box has been improved.

AO control SW

1) Development of test and calibration GUI is in progress.
2) implementation of high order force filtering.

LBT672a

1) The 4D interferometer has been mounted into the tree house and checked remote desktopping.
2) The switch on test of the interferometer failed. test of the interferometer connections is planned for tomorrow.

Meeting: A meeting of all the Arcetri group was done today to decide about AO commissioning schedule for the period April-June. The travel plans of all the group was arranged to fit with the schedule.

Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback